Home‎ > ‎Reports‎ > ‎Crew 108‎ > ‎

day07


Crew #108 Log Book for December 09, 2011

Photos of the day





Commander's Check-In Report

Date and time: 

9 December 2011


Written by : 

Charlotte Poupon, Crew 108, Commander


Crew Physical Status:

The 5 crew members remaining are in a gorgeous state of health and mood


Time Departed/Returned from EVA:

EVA # 13 / 10:12 to 12:20
EVA # 14 / 10:48 to 11:59


Brief Narrative of Field Mission Results:

Astro-biologist took samples as she planned to.


EVA Data/Interpretations:

The second EVA settled a meeting point down to Phobos Peek to climb it again as it seems to be a nice spot for sampling.

Engineering/Hab Maintenance:

It is the first day that Engineers did not work on something specific, since everything is working well.


Report Transmission Schedule:

All reports by 22:00.


Plans for Tomorrow:

This is our mid-rotation. Will make it like an amazing day = lunar eclipse, then scrambled eggs and bacon...


Inventory:

I deeply suggest the MDRS to buy a converter 110/220 v since the Rover we plan to operate is stuck because of its unfit plug (and also its conceptor - the commander - that missed that point)


Support Requested:

If any body has this adaptator it can be great... if not... ok...


Miscellaneous:

Commander's Report

Date and time: 

9 December 2011

Written by :

Charlotte Poupon, Crew 108, Commander


Text:

This days was nice and pleasant at the MDRS. All crew members fit perfectly 
together, duties of the simulation and the all constraints on life.

Engineers have a small break from the past few days, since everything is operational in the Hab due to their efficiency of the past days. Tonight we'll go to sleep maybe sooner than other days to wake up around 5am for the lunar eclipse.

Science Report

Date and Time: 

19:07, 9th December, 2011


Written by: 

Usha Lingappa, Crew 108, Astrobiologist

On both EVAs today, I took samples of rock varnish; first from the summit of Phobos Peak, and then from the dinosaur excavation site at the end of Lowell Highway. I brought it back to the lab and surface sterilized the rocks with ethanol, and then aseptically scraped the varnish from the rocks using a metal file under the flame of a blowtorch. I made a slurry of the powdered varnish in sterile ultra-pure water, and inoculated two Luria Bertani broth agar plates with varnish slurry which I will incubate overnight.


Engineering Report

Date and time: 

9 December 2011


Written by: 

Mike Lotto, Crew 108, Engineer


Generator/Electricity:

Kitty (Diesel Generator) (Used? Yes/No): Yes
Notes/Comments (include how many generator runs, approximately when,
and any times you needed to turn the generator on/off):

Oil:

Kitty Oil Quantity (# of Quarts in storage): Unknown
Honey Oil Quantity (# of Quarts in storage): Unknown
ATV Oil Quantity (# of Quarts in storage): 3
Car Oil Quantity (# of Quarts in storage): 3

Water Status:

Outside Potable Water Tank Level (inches from bottom): 0.5
Trailer Potable Water Tank Level (inches from bottom): 14
Hab Potable Water Tank Level (inches from bottom): 12

Observatory:

Telescope Used? (Yes/No): No
Observatory Notes/Comments: N/A

Transportation:

Hab Car (NOT RENTAL)
Used (yes/no): Yes
Oil Checked (yes/no):No
Tire Status: Good
Notes/Comments on Hab Car: N/A

ATV 1:
Used: Yes
Oil Checked: No
Fuel Consumed: 0.2 gallons
Tire Status: Good
Comments: 7 remaining gallons of spare fuel for all ATVs

ATV 2:
Used: Yes
Oil Checked: No
Fuel Consumed: 0.2 gallons
Tire Status: Good
Comments: 7 remaining gallons of spare fuel for all ATVs

ATV 3:
Used: Yes
Oil Checked: No
Fuel Consumed: 0.2 gallons
Tire Status: Good
Comments: 7 remaining gallons of spare fuel for all ATVs

Heating and Ventilation:

Thermostat setting upstairs (Farenheit): 65
Actual temperature upstairs (Farenheit): 69.4
Thermostat setting downstairs (Farenheit): 55
Actual temperature downstairs (Farenheit): 50

Computers/Networking Infrastructure:

Notes/comments on internet/computers?: 

Regarding the VisionGS, the http port is set to 80 - but nothing is uploading


General Engineering Notes/Comments: 

Will look into securing joints on EVA Suit 5

Journalist Report

Date and Time: 

18:17, 9th December, 2011


Written by: 

Usha Lingappa, Crew 108, Astrobiologist


Today's main events were two full-crew EVAs. In the morning, our destination was Phobos Peak (where EVA 11 went yesterday). Two crew members, Lotto and LeClair, set out by foot about 30 minutes before Dale, Poupon, and I followed by ATV. We rendezvoused at the base of Phobos Peak, and together climbed to the summit. 

I took some rock varnish samples while up there. We all returned to the Hab by lunch time. In the late afternoon, we all piled into the pressurized rover for a second EVA. We drove out to the end of Lowell Highway to the dinosaur excavation to try and play with Poupon's small rover, whom we call BuddyBot. 

Unfortunately, due to incompatibility between European and North American power supplies BuddyBot is not working for the time being. The drive back to the Hab was stunningly beautiful. Although on some level this completely broke the simulation, we watched a spectacular moonrise. We are looking forward to the lunar eclipse early tomorrow morning.


EVA 13 Report

Date and time: 

9 December 2011


Written by:

Michael LeClair, Crew 108, XO


Time: 

10:12 - 12:20


Crew Members: 

LeClair, Lotto


Site Location: 

Phobos Peak


Transit Mode:

Foot travel


Objectives: 

Multi-crew field rendezvous, geological sample collection, photography


Results: 

Multi-crew field rendezvous maneuvers pose no challenges. 7 cm chalcedony spheroid sample collected. Photos taken at Mellow Mushrooms and of tripped Dragon from the North. Photos taken from peak in all directions.


Lessons Learned: 

Plains are a place of geological interest. Sense of scale difficult for longer distances in the absence of distinguishing features during foot traverses.


EVA 14 Report

Date and time: 

9 December 2011


Written by: 

Ashley Dale, Crew 108, Engineer


Time: 

10:43 – 11:59

- Depressurization: 10:43 – 10:48

- ATV to base of Phoebes Peak: 10:48 – 10:53

- Walk from ATV to Rendezvous point: 10:53 – 11:00

- Rendezvous with Mike Lotto (Engineer) and Michael LeClair (Geologist): 11:00

- Ascend Phoebes Peak: 11:02 – 11:17

- Loiter at top of Phoebes Peak and collect rock varnish: 11:17 – 11:33

- Descend Phoebes Peak and return to ATVs: 11:33 – 11:46

- ATV back to Hab: 11:46 – 11:53

- Repressurization: 11:54 - 11:59

Crew Members:

Charlotte Poupon (Station Leader), Usha Lingappa (Astrobiologist), Ashley Dale (Engineer)


Site Location: 

Phobos Peak


Transit Mode: 

ATV 1-3

Objectives: 

Rendezvous with Mike Lotto (Engineer) and Michael LeClair (Geologist) at base of Phoebes Peak, collect rock varnish from top of Phobos Peak.


Results: 

Rendezvous completed successfully, top of Phobos peak reached, rock varnished attained for further study.


Lessons Learned: 

Rock hammers are difficult to use (Lingappa), the plains to the west of Phobos Peak are of geological interest, temperature change over the morning is quite rapid (other EVA team recommended dressing warm, 20 minutes later it was warm enough to dress lightly).

Chef's Report

Prepared by Charlotte Poupon, Mike Lotto of Crew 108

1a) Today was
- a cooking day

1b) Was today a special day celebrated at the main meal with special
food or activity? No

2) List the foods served at today's main meal (usually the evening
meal), giving full names of each.

We cooked a risotto. We had rice + deh
ydrated onions + green beans + crushed nuts.

We miss a tiny little simple glass of white wine to start the rice cooking !

3) List any main meal foods not finished by the crew (leftovers stored
for later use or discarded)

1 serving of risotto.

We also through away the try we did with yagourth since it did not
succeed AT ALL... Are we sure it works will dry milk ?

4) Recipes for anything more complicated than following package
directions exactly, or rehydrating.

- boil 1,5 liter of water with vegetable stock
- re-hydrate the onions
- re-hydrate the green beans
- crush nuts

- fried the onions on butter+olive oil.
- Add the rice.
- Add little by little the stock and wait the rice absorbe it before adding more.
- leave 10 min with the lid on.
- add the green beans and nuts.
- mixed, salt + pepper...

HERE IT IS :)


5) Lessons learned (if any)

O


6) Comments/questions for Kim and Jean, the food study investigators

- Should be nice to make a real inventory of all spices, and buy new because it can 
change the recipes so much !!!

- We missed lemon juice, so may be the kind of concentrate one can be applicable for food study.

- 1 tortilla/day is somehow not enough, since we can enjoy it for breakfast +meal (or snack)... 2 would be better...